OpenStudy (anonymous):

Solve each system graphically. Be sure to check your solution. If a system has an infinite number of solutions, use set−builder notation to write the solution set. If a system has no solution, state this. 6x-2y=2 9x-3y=1

OpenStudy (anonymous):

no solutions

OpenStudy (anonymous):

the row reduced echelon form of {{6,-2 2},{9,-3,1}} is inconsistent.

OpenStudy (anonymous):

Or the determinant of the coefficient matrix = 0, and divisibility by zero is undefined so there are no solutions for either x of y therefore the system of equations is inconsistent

OpenStudy (anonymous):

Or the graphs of 6x-2y=2 and 9x-3y=1 are lines with no intersection (parallel to each other) 6x-2y=2 can also be written as 3x-1=y 9x-3y=1 can also be written as 3x-1/3=y both a parallel to each other